# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix H Extension to Continuous State/Action Spaces
Definition H.10.
Let the complete feature map be denoted by $\Phi\in\mathbb{R}^{|\mathcal{S}|\times d}$, where
$\Phi=\begin{bmatrix}\phi(1)\\ \phi(2)\\ \vdots\\ \phi(S)\end{bmatrix}\in\mathbb{R}^{S\times d}$ (51)
Similarly, let the subsampled feature map be denoted by $\widehat{\Phi}_{k}\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|\times d}$, where
$\widehat{\Phi}_{k}=\begin{bmatrix}\bar{\phi}_{k}(1)\\ \bar{\phi}_{k}(2)\\ \vdots\\ \bar{\phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|)\end{bmatrix}\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|\times d}\$ (52)
Here, the system’s stage rewards are given by :
$r=\begin{bmatrix}r(1)&\dots&r(S)\end{bmatrix}\in\mathbb{R}^{S},$ (53)
$r_{k}=[r_{k}(1),\dots,r_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|]\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|},$ (54)
where $d$ is a low-dimensional embedding we wish to learn.
@@
$V_{w}\approx\widehat{\Phi}_{k}w=\begin{bmatrix}\widehat{\Phi}_{k}(1)\\ \vdots\\ \widehat{\Phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})})\end{bmatrix}w\in\mathrm{span}(\widehat{\Phi}_{k})$
This manner of updating the weights can be considered via the projected Bellman equation $\widehat{\Phi}_{k}w=\Pi_{\mu}\mathcal{T}(\widehat{\Phi}_{k}w)$, where $\Pi_{\mu}(v)=\arg\min_{z\in\widehat{\Phi}_{k}w}\|z-v\|_{\mu}^{2}$. Notably, the fixed point of the the projected Bellman equation satisfies $w=(\widehat{\Phi}_{k}^{\top}D_{\mu}\widehat{\Phi}_{k})^{-1}\widehat{\Phi}_{k}^{\top}D_{\mu}(r+\gamma P^{\pi}\widehat{\Phi}_{k}w)$, where $D_{\mu}$ is a diagonal matrix comprising of $\mu$’s. In other words, $D_{\mu}=\mathrm{diag}(\mu_{1},\dots,\mu_{n})$. Then,
$(\widehat{\Phi}_{k}^{\top}D_{\mu}\widehat{\Phi}_{k})w=\widehat{\Phi}_{k}^{\top}D_{\mu}(r+\gamma P^{\pi}\widehat{\Phi}_{k}w)$
@@
$\widehat{\Phi}_{k}^{\top}D_{\mu}(I-\gamma P^{\pi})\Phi w=\widehat{\Phi}_{k}^{\top}D_{\mu}r.$
Therefore, the problem is amenable to Algorithm 1 in Ren et al. (2024). To bound the error of using linear function approximation to learn $\widehat{Q}_{k}$, we use a result from Ren et al. (2024).
Lemma H.11 (Policy Evaluation Error, Theorem 6 of Ren et al. (2024)).
Suppose the sample size $M\geq\log\left(\frac{4n}{\delta}\right)$, where $n$ is the number of agents and $\delta\in(0,1)$ is an error parameter. Then, with probability at least $1-2\delta$, the ground truth $\widehat{Q}^{\pi}_{k}(s,a)$ function and the approximated $\widehat{Q}_{k}$-function $\widehat{Q}^{\mathrm{LFA}}_{k}(s,a)$ satisfies, for any $(s,a)\in\mathcal{S}\times\mathcal{A},$
@@
Therefore, when ${\pi}_{k,m}$ is derived from the spectral features learned in $\widehat{Q}^{\mathrm{LFA}}_{k}$, applying the triangle inequality on the Bellman noise and setting $\delta=\epsilon_{P}=1/\sqrt{k}$ yields [^1] Following Ren et al. (2024), the result easily generalizes to any positive-definite transition Kernel noise (e.g. Laplacian, Cauchy, Matérn, etc.
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}+\log\left(2k^{2}\right)\frac{\|\bar{\phi}_{k}\|^{5}}{\sqrt{M}}+\frac{2}{\sqrt{k}}\cdot\frac{\gamma\widetilde{r}}{1-\gamma}\|\bar{\phi}_{k}\|\right)\right]\geq\left(1-\frac{1}{100\sqrt{k}}\right)\left(1-\frac{2}{\sqrt{k}}\right)$ (55)
Using $\|\widetilde{\phi}_{k}\|\leq 1$ gives
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}+\frac{\log\left(2k^{2}\right)}{\sqrt{M}}+\frac{2\gamma\widetilde{r}}{\sqrt{k}(1-\gamma)}\right)\right]\geq\left(1-\frac{1}{100\sqrt{k}}\right)\left(1-\frac{2}{\sqrt{k}}\right)$
Remark H.12.
@@
Hence, in the continuous state/action space setting, as $k\to n$ and $M\to\infty$, $V^{\widetilde{\pi}_{k}}(s_{0})\to V^{\widetilde{\pi}_{n}}(s_{0})=V^{\pi^{*}}(s_{0})$. Intuitively, as $k\to n$, the optimality gap diminishes following the arguments in Corollary F.6.1. As $M\to\infty$, the number of samples obtained allows for more effective learning of the spectral features.
Appendix I Towards Deterministic Algorithms: Sharing Randomness
In this section, we propose algorithms RANDOM-SUB-SAMPLE-MFQ and RANDOM-SUB-SAMPLE-MFQ+, which shares randomness between agents through various careful groupings and using shared randomness within each groups. By implementing these algorithms, we derive simulation results, and establish progress towards derandomizing inherently stochastic approximately-optimal policies.
Algorithm 6 (Execution with weakly shared randomness). The local agents are heuristically divided into arbitrary groups of size $k$. For each group $h_{i}$, the $k-1$ local agents sampled are the same at each iteration. The global agent’s action is then the majority global agent action determined by each group of local agents. At each round, this requires $O(\left\lceil n^{2}/k\right\rceil(n-k))$ random numbers.
Algorithm 6 SUB-SAMPLE-MFQ: Execution with weakly shared randomness
0:  A multi-agent system as described in Section 2. A distribution $s_{0}$ on the initial global state $s_{0}=(s_{g},s_{{[n]}})$. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})$.
1:  Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$.
2:  Define groups $h_{1},\dots,h_{x}$ of agents where $x\coloneq\left\lceil\frac{n}{k}\right\rceil$ and $|h_{1}|=|h_{2}|=\dots=|h_{x-1}|=k$ and $|h_{x}|=n\mod k$.
3:  for $t=0$ to $T^{\prime}-1$ do
4:     for $i\in[x]$ do
5:        Let $\Delta_{i}$ be a uniform sample of (${[n]\setminus g_{i}\choose k-1}$, and let $a_{g_{i}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta_{i}}(t))]_{1:k}$.
6:     Let $a_{g}(t)=\mathrm{majority}\left(\left\{[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta_{i}}(t))]_{g}\right\}_{i\in[x]}\right)$.
7:     Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t))$
8:     Let $s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t)$
9:     $s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
10:     Get reward $r(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g})$
Algorithm 7 (Execution with strongly shared randomness). The local agents are randomly divided in to groups of size $k$. For each group, each agent uses the other agents in the group as the $k-1$ other sampled states. Similarly, the global agent’s action is then the majority global agent action determined by each group of local agents. Here, at each round, this requires $O(\left\lceil n^{2}/k\right\rceil)$ random numbers.
Algorithm 7 SUB-SAMPLE-MFQ: Execution with strongly shared randomness
0:  A multi-agent system as described in Section 2. A distribution $s_{0}$ on the initial global state $s_{0}=(s_{g},s_{{[n]}})$. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{\Delta})$.
1:  Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$.
2:  Define groups $h_{1},\dots,h_{x}$ of agents where $x\coloneq\left\lceil\frac{n}{k}\right\rceil$ and $|h_{1}|=|h_{2}|=\dots=|h_{x-1}|=k$ and $|h_{x}|=n\mod k$.
3:  for $t=0$ to $T^{\prime}-1$ do
4:     for $i\in[x-1]$ do
5:        Let $a_{h_{i}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{i}}(t))]_{1:k}$.
6:        Let $a^{g}_{h_{i}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{i}}(t))]_{g}$.
7:     Let $\Delta_{\mathrm{residual}}\coloneq{[n]\setminus h_{x}\choose k-(n-\left\lceil n/k\right\rceil k)}$.
8:     Let $a_{h_{x}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{x}\cup\Delta_{\mathrm{residual}}}(t))]_{1:n-\left\lceil n/k\right\rceil k}$
9:     Let $a^{g}_{h_{x}}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{h_{x}\cup\Delta_{\mathrm{residual}}}(t))]_{g}$
10:     Let $a_{g}(t)=\mathrm{majority}\left(\{a_{h_{i}}^{g}(t):i\in[x]\}\right)$.
11:     Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t)),s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
12:     Get reward $r(s,a)=r_{g}(s_{g},a_{g})+\frac{1}{n}\sum_{i\in[n]}r_{l}(s_{i},a_{i},s_{g})$
The probability of a bad event (the policy causing the $Q$-function to exceed the Lipschitz bound) scales with the $O(n^{k})$ independent random variables, which is polynomially large. Agnostically, some randomness is always needed to avoid periodic dynamics that may occur when the underlying Markov chain is not irreducible. In this case, an adversarial reward function could be designed such that the system does not minimize the optimality gap, thereby penalizing excessive determinism.
This ties into an open problem of interest, which has been recently explored (Larsen et al., 2024). What is the minimum amount of randomness required for SUBSAMPLE-MFQ (or general mean-field or multi-agent RL algorithms) to perform well? Can we derive a theoretical result that demonstrates and balances the tradeoff between the amount of random bits and the performance of the subsampled policy when using the greedy action from the $k$-agent subsystem derived from $\widehat{Q}_{k}^{*}$? We leave this problem as an exciting direction for future research.
Appendix J Numerical Experiments
This section provides simulation results for the examples outlined in Section 2 for Gaussian squeeze and constrained exploration. All numerical experiments were run on a 3-core CPU server equipped with a RAM. We chose a complexity for the parameters of each simulation that was sufficient to emphasize characteristics of the theory, such as the complexity improvement of Algorithms 1 and 2 and the decaying optimality gap.
J.1 Gaussian Squeeze
In this task, $n$ homogeneous agents determine their individual action $a_{i}$ to jointly maximize the objective $r(x)=xe^{-(x-\mu)^{2}/\sigma^{2}}$, where $x=\sum_{i=1}^{n}a_{i}$, $a_{i}=\{0,\dots,9\}$, and $\mu$ and $\sigma$ are the pre-defined mean and variance of the system. In scenarios of traffic congestion, each agent $i\in[n]$ is a traffic controller trying to send $a_{i}$ vehicles into the main road, where controllers coordinate with each other to avoid congestion, hence avoiding either over-use or under-use, thereby contributing to the entire system. This GS problem serves as an ablation study on the impact of subsampling for MARL.
Specifically, we consider a purely homogeneous subsystem devoid of a global agent to match the mean-field MARL setting in Yang et al. (2018), where the richness of our model setting can still express the Gaussian squeeze task. Instead, we use the global agent to model the global state of the system, given by the number of vehicles in each controller’s lane.
We set $\mathcal{S}_{l}=\{1,\dots,20\}$, $\mathcal{A}_{l}=\{0,\dots,9\}$, and $\mathcal{S}_{g}=\mathcal{S}_{l}$ such that $s_{g}=\left\lceil\frac{1}{n}\sum_{i=1}^{n}s_{i}\right\rceil$. The transition functions are given by $s_{i}(t+1)=s_{i}(t)-1\cdot\mathbbm{1}\{s_{i}(t)>s_{g}(t)\}+\mathrm{Ber}(p)$ and $s_{g}(t+1)=\left\lceil\frac{1}{n}\sum_{i=1}^{n}s_{i}(t+1)\right\rceil$, where $\mathrm{Ber}(p)$ is a Bernoulli random variable with parameter $p>0$. Finally, the global agent’s reward function is given by $r_{g}(s_{g})=-s_{g}$ and the local agent’s reward function is given by $r_{l}(s_{i},s_{g},a_{i})=4\cdot\mathbbm{1}\{s_{i}>s_{g}\}-2\cdot\mathbbm{1}\{a_{i}>s_{g}\}$. For this task, we ran two simulations: a large-scale simulation with $n=50$ agents, and a small-scale simulation with $n=6$ agents, and provide simulation results below. We observe the exponential improvement in computational complexity in Figure 6, as well as the monotonic improvement in the cumulative reward in Figure 7.
Figure 6: Computation time (in minutes) against sampling parameter $k$.
Figure 7: Cumulative discounted reward as $k$ increases
J.2 Constrained Exploration
Consider an $M\times M$ grid. Each agent’s state is a coordinate in $[M]\times[M]$. The state represents the center of a $d\times d$ box where the global agent wishes to constrain the local agents’ movements. Initially, all agents are in the same location. At each time-step, the local agents take actions $a_{i}(t)\in\mathbb{R}^{2}$ (e.g., up, down, left, right) to transition between states and collect stage rewards. The transition kernel ensures that local agents remain within the $d\times d$ box dictated by the global agent, by only using knowledge of $a_{i}(t),s_{g}(t)$, and $s_{i}(t)$. In warehouse settings where some shelves have collapsed, creating hazardous or inaccessible areas, we want agents to clean these areas. However, exploration in these regions may be challenging due to physical constraints or safety concerns. Therefore, through an appropriate design of the reward and transition functions, the global agent could guide the local agents to focus on specific $d\times d$ grids, allowing efficient cleanup while avoiding unnecessary risk or inefficiency.
We set $\mathcal{S}_{g}=\mathcal{S}_{l}=\{0,1,2,3,4,5,6\}\times\{0,1,2,3,4,5,6\}$ and $\mathcal{A}_{l}=\mathcal{A}_{g}=\left\{\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\end{pmatrix},\begin{pmatrix}1\\ 0\end{pmatrix},\begin{pmatrix}1\\ 1\end{pmatrix}\right\}$. Then:
$s_{g}^{t+1}(s_{g}^{t},a_{g}^{t})=\Pi_{S_{g}}(s_{g}^{t}+a_{g}^{t})$
$s_{i}^{t+1}(s_{i}^{t},s_{g}^{t},a_{i}^{t})=\Pi_{S_{l}}(s_{i}^{t}+|s_{i}^{t}-s_{g}^{t}|+a_{i}^{t})$
$r_{g}(s_{g}^{t},a_{g}^{t})=10*\mathbbm{1}\{a_{g}^{t}=0\}$
$r_{l}(s_{i}^{t},s_{g}^{t},a_{i}^{t})=\mathbbm{1}\{a_{i}^{t}=0\}+4\cdot\mathbbm{1}\{|s_{i}^{t}-s_{g}^{t}|=1\}-2\cdot|s_{i}^{t}-s_{g}^{t}|$
For this task, we ran a simulation with $n=20$ agents, and provide simulation results in Fig. 8. Again, we observe the monotonic improvements in the cumulative discounted rewards.
Figure 8: Reward optimality gap (log scale) with ${\pi}_{k,m}^{\mathrm{est}}$ running $300$ iterations.

### v3

@@ Appendix I Partially Relaxing the Offline Learning assumption: Off-Policy Learning
Definition J.11.
Let the complete feature map be denoted by $\Phi:\mathcal{S}\times d\to\mathbb{R}$, and let the subsampled feature map be denoted by $\hat{\Phi}_{k}:\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times d\in\mathbb{R}$, where we let
$\hat{\Phi}_{k}=\begin{bmatrix}\bar{\Phi}_{k}(1)\\ \bar{\Phi}_{k}(2)\\ \vdots\\ \bar{\Phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|)\end{bmatrix}\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|\times d}\$ (44)
Here, the system’s stage rewards are given by $r=\begin{bmatrix}r(1)&\dots&r(S)\end{bmatrix}\in\mathbb{R}^{S}$ and $r_{k}=[r_{k}(1),\dots,r_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|]\in\mathbb{R}^{|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})}|}$, where $d$ is a low-dimensional embedding we wish to learn.
@@
$V_{w}\approx\hat{\Phi}_{k}w=\begin{bmatrix}\hat{\Phi}_{k}(1)\\ \vdots\\ \hat{\Phi}_{k}(|\mathcal{S}_{g}|\times|\mu_{k-1}{(\mathcal{S}_{l})})\end{bmatrix}w\in\mathrm{span}(\hat{\Phi}_{k})$
This manner of updating the weights can be considered via the projected Bellman equation $\hat{\Phi}_{k}w=\Pi_{\mu}\mathcal{T}(\hat{\Phi}_{k}w)$, where $\Pi_{\mu}(v)=\arg\min_{z\in\hat{\Phi}_{k}w}\|z-v\|_{\mu}^{2}$. Notably, the fixed point of the the projected Bellman equation satisfies
$w=(\hat{\Phi}_{k}^{\top}D_{\mu}\hat{\Phi}_{k})^{-1}\hat{\Phi}_{k}^{\top}D_{\mu}(r+\gamma P^{\pi}\hat{\Phi}_{k}w),$
where $D_{\mu}$ is a diagonal matrix comprising of $\mu$’s. In other words, $D_{\mu}=\mathrm{diag}(\mu_{1},\dots,\mu_{n})$. Then,
$(\hat{\Phi}_{k}^{\top}D_{\mu}\hat{\Phi}_{k})w=\hat{\Phi}_{k}^{\top}D_{\mu}(r+\gamma P^{\pi}\hat{\Phi}_{k}w)$
@@
$\hat{\Phi}_{k}^{\top}D_{\mu}(I-\gamma P^{\pi})\Phi w=\hat{\Phi}_{k}^{\top}D_{\mu}r.$
Therefore, the problem is amenable to Algorithm 1 in Ren et al. [2024]. To bound the error of using linear function approximation to learn $\hat{Q}_{k}$, we use a result from Ren et al. [2024].
Lemma J.12 (Policy Evaluation Error, Theorem 6 of Ren et al. [2024]).
Suppose the sample size $M\geq\log\left(\frac{4n}{\delta}\right)$, where $n$ is the number of agents and $\delta\in(0,1)$ is an error parameter. Then, with probability at least $1-2\delta$, the ground truth $\hat{Q}^{\pi}_{k}(s,a)$ function and the approximated $\hat{Q}_{k}$-function $\hat{Q}^{\mathrm{LFA}}_{k}(s,a)$ satisfies, for any $(s,a)\in\mathcal{S}\times\mathcal{A},$
@@
Therefore, when ${\pi}_{k,m}$ is derived from the spectral features learned in $\hat{Q}^{\mathrm{LFA}}_{k}$, applying the triangle inequality on the Bellman noise and setting $\delta=\epsilon_{P}=\frac{1}{\sqrt{k}}$ yields [^3] Following Ren et al. [2024], the result easily generalizes to any positive-definite transition Kernel noise (e.g. Laplacian, Cauchy, Matérn, etc.
$\Pr\left[V^{\pi^{*}}(s_{0})\!-\!V^{{\pi}_{k,m}}(s_{0})\!\leq\!\tilde{O}\left(\frac{1}{\sqrt{k}}+\log\left(2k^{2}\right)\frac{\|\bar{\phi}_{k}\|^{5}}{\sqrt{M}}\!+\!\frac{2}{\sqrt{k}}\cdot\frac{\gamma\tilde{r}}{1-\gamma}\|\bar{\phi}_{k}\|\right)\right]\!\geq\!1\!+\!\frac{1}{50k}\!-\!\frac{201}{100\sqrt{k}}$
Using $\|\tilde{\phi}_{k}\|\leq 1$ and simplifying gives
$\Pr\left[V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\tilde{O}\left(\frac{1}{\sqrt{k}}+\frac{\log\left(2k^{2}\right)}{\sqrt{M}}+\frac{2\gamma\tilde{r}}{\sqrt{k}(1-\gamma)}\right)\right]\geq 1-\frac{201}{100\sqrt{k}}$
Remark J.14.
@@
Hence, in the continuous state/action space setting, as $k\to n$ and $M\to\infty$, $V^{\tilde{\pi}_{k}}(s_{0})\to V^{\tilde{\pi}_{n}}(s_{0})=V^{\pi^{*}}(s_{0})$. Intuitively, as $k\to n$, the optimality gap diminishes following the arguments in Corollary G.7. As $M\to\infty$, the number of samples obtained allows for more effective learning of the spectral features.